# DCF-DS: Deep Cascade Fusion of Diarization and Separation for Speech Recognition under Realistic Single-Channel Conditions

### v1

@@ Abstract
We propose a single-channel Deep Cascade Fusion of Diarization and Separation (DCF-DS) framework for back-end speech recognition, combining neural speaker diarization (NSD) and speech separation (SS). First, we sequentially integrate the NSD and SS modules within a joint training framework, enabling the separation module to leverage speaker time boundaries from the diarization module effectively. Then, to complement DCF-DS training, we introduce a window-level decoding scheme that allows the DCF-DS framework to handle the sparse data convergence instability (SDCI) problem. We also explore using an NSD system trained on real datasets to provide more accurate speaker boundaries during decoding. Additionally, we incorporate an optional multi-input multi-output speech enhancement module (MIMO-SE) within the DCF-DS framework, which offers further performance gains. Finally, we enhance diarization results by re-clustering DCF-DS outputs, improving ASR accuracy. By incorporating the DCF-DS method, we achieved first place in the realistic single-channel track of the CHiME-8 NOTSOFAR-1 challenge. We also perform the evaluation on the open LibriCSS dataset, achieving a new state-of-the-art performance on single-channel speech recognition.
@@
Multi-speaker speech recognition aims to identify “who spoke what and when” in recordings with unknown scenarios and speakers [1]. While deep learning has driven significant advancements in speech recognition [2, 3], accurate speech recognition in real-world multi-speaker conversational scenarios remains a major challenge. These challenges often arise from high overlap rates, background noises and reverberation, unknown speaker numbers, and natural conversation styles [4, 5, 6]. In this context, speaker diarization and speech separation are commonly employed as pre-processing steps. Specifically, speaker diarization segments the audio based on speaker identity [7, 8], while speech separation [9] isolates each speaker’s speech from the mixed audio signals.
@@
Speaker diarization and speech separation are closely related [4], as both aim to predict the distribution of speakers: speaker diarization predicts the speaker’s presence across time frames, whereas speech separation predicts their distribution across time-frequency (T-F) bins. Consequently, the temporal distribution provided by diarization can serve as a prior for speech separation, helping the system to better distinguish between speakers [10, 11, 12, 13, 4, 14]. Speech separation, in turn, inherently encompasses diarization—ideal separation results delineate the time regions of speakers [15, 16, 17]. Moreover, speech separation can handle overlapping segments, which can enhance speaker diarization accuracy [18, 6]. Due to the strong correlation between speaker diarization and speech separation, numerous kinds of research [19, 20, 21, 18, 6, 15, 16, 17, 10, 11, 12, 13, 4, 14, 22] have been conducted to explore combining these two tasks to enhance speech quality. Based on how the two tasks are combined, these methods can be broadly categorized into multi-task joint training and sequential cascade approaches.
@@
In multi-task joint training approaches [19, 20, 22, 21], the network typically generates two outputs: speaker activity probabilities and separated speech signals. This structure allows the two tasks to benefit from each other mutually [20]. Speaker diarization results can also be used to post-process the separation outputs, reducing background noises in the separated signals [19, 20, 21, 22]. Among these methods, some studies [19, 22] propose a network that jointly performs target speech extraction and voice activity detection (VAD), yielding improvements in both fully and sparsely overlapped speech. Another approach, end-to-end neural speaker diarization and separation (EEND-SS) [20], integrates speaker counting, diarization, and separation in a multi-tasking fashion. However, due to the inclusion of the separation task within the joint training framework, these methods are restricted to training on simulated data. Unsupervised training methods, such as PixIT [21], can help mitigate this issue. However, when the number of separated output nodes is limited, clustering methods are still required to assign separated streams for long recordings.
@@
In sequential cascade framework [18, 6, 15, 16, 17, 10, 11, 12, 13, 4, 14], there is no definitive answer as to whether speech separation or speaker diarization should be performed first. Some approaches [18, 6, 15, 16, 17] choose to perform speech separation first, namely “separation-then-diarization”. For example, the continuous speaker separation (CSS) pipeline [18, 6] first processes overlapping regions in the original audios through speech separation and then applies the separated results to diarization. Some studies [15, 16, 17] explore speaker diarization by performing VAD on the separated outputs, such as speech separation guided diarization (SSGD) [15] and recurrent selective attention network (RSAN) [16, 17]. Some methods [10, 11, 12, 13, 4, 14] begin with speaker diarization and then use the obtained boundary information to guide speech separation, namely “diarization-then-separation”. This is also a promising direction, as speaker diarization is generally simpler than speech separation, allowing the former to provide a more accurate prior. A classic method is guided source separation (GSS) [10, 11], which uses segment boundaries to initialize time-varying mixture weights based on traditional spatial mixture models [23, 24]. However, GSS relies on multi-channel spatial information. The speaker activity driven speech extraction neural network (ADEnet) [25] explores how to use speaker activity boundaries as auxiliary cues for speech extraction. The CSS-AD [26] pipeline performs speech recognition on CSS outputs, followed by diarization on the ASR output, achieving promising results on the LibriCSS dataset [18]. Certain approaches [12, 13] enhance separation performance in real two-speaker scenarios by using speaker diarization to generate adaption training data. Recently, the target-speaker separation (TS-SEP) [4] and speaker separation via neural diarization (SSND) [14] have achieved strong performance in the fields of diarization and separation. TS-SEP method [4] combines speaker diarization and speech separation by extending the output of target-speaker voice activity detection (TS-VAD) [27] from the time domain to the time-frequency domain. SSND [14] uses estimated speaker boundaries to assist in assigning speakers to the outputs of a multi-speaker separation model. These two methods demonstrate the great potential of the sequential cascade framework.
@@
From various studies mentioned above, it is apparent that both joint training and sequential cascade approaches have advantages and limitations. Joint training allows simultaneous diarization and separation within a compact and unified framework. However, typical multi-task learning frameworks usually do not fully leverage diarization results to optimize separation results. Additionally, joint training frameworks are generally restricted to simulated data. Conversely, the sequential cascade approach—particularly “diarization-then-separation”—allows for more effective integration of two tasks, where separation can effectively leverage diarization results. However, this approach tends to be with a more cumbersome process and independently optimized components. Therefore, an effective approach is to combine “joint training” and “sequential cascade“, thereby taking advantage of both methods—what we refer to as “deep cascade fusion”. However, due to differences in diarization and separation tasks, direct integration introduces a problem: diarization models can have many output nodes (e.g., eight output nodes in [28]). Directly feeding these outputs into the separation model and aligning the number of separation outputs with those from diarization may lead to non-convergence [4]. This issue arises because the separation model usually allocates an output node for each speaker. However, when training on typical conversational data, each node has limited target signals, making it difficult for the neural network to learn speaker-discriminative features. For example, in a natural eight-speaker meeting conversation, there are many regions where only one or two people are speaking, a separation model with eight outputs will have only one or two nodes with effective training targets in those regions. This significantly increases the training difficulty, which we term the sparse data convergence instability (SDCI) problem. This issue was also highlighted when the TS-SEP method was proposed [4], where a two-stage training approach was proposed as the solution. However, two-stage training makes the TS-SEP method unable to fully leverage the boundary information from the diarization module for separation. The SSND addresses this issue by constructing dual embedding sequences, which leads to a complex framework. This approach requires a pre-trained diarization model to extract speaker embeddings, followed by constructing an embedding sequence based on boundary information during both the training and testing phases. Additionally, a scenario-aware differentiated loss [22, 29] has been proposed to handle the sparsely overlapped speech. However, this approach requires defining multiple different loss functions and prior knowledge of each speaker’s VAD information.
@@
To address the above issues, we propose a single-channel deep cascade fusion of diarization and separation (DCF-DS) framework for back-end multi-speaker speech recognition. The primary contributions of our study are outlined as follows:
@@
In DCF-DS, a joint training framework sequentially combining neural speaker diarization and separation is designed. Compared to TS-SEP, DCF-DS effectively leverages temporal boundaries from the diarization, thereby reducing the difficulty of separation. Compared to SSND, DCF-DS avoids complex pre-processes and eliminates dependence on spatial information. To complement DCF-DS training, we design a window-level decoding scheme to handle the SDCI problem. This decoding scheme allows the separation module to produce fewer outputs than the total number of speakers in an utterance.
@@
The remainder of this paper is organized as follows. Section II provides an overview of related prior works. Section III describes the proposed DCF-DS framework in detail. In Section IV, we present and analyze the experimental results. Finally, Section V concludes the paper and discusses potential directions for future research.
@@ II-A Neural Speaker Diarization
where represents the input acoustic features. In NSD-MS2S, we use log-Mel filter-bank features (FBANKs). denotes the -dimensional feature vector at frame , where represents the total number of frames. represents the speaker embedding vector of the -th speaker, and denotes the number of speakers. is the speaker mask matrix, which has elements that indicate whether speaker is present at frame , taking a value of 1 if present and 0 if not. The speaker mask matrix is obtained through manual annotations during training and an initialized diarization system (e.g., spectral clustering) during testing. The NSD-MS2S transforms the inputs into deep features through the following process:
where denotes the convolutional layers, is the number of output channels. is the new feature dimension after downsampling. represents the memory-aware multi-speaker embedding (MA-MSE) module proposed in [42], which can provide dynamically refined speaker embeddings using the speaker embedding basis matrix . Next, NSD-MS2S passes these features through an encoder-decoder architecture to generate the final output:
@@ II-B Continuous Speech Separation
where denotes the T-F mask for speaker . denotes the number of outputs. In traditional CSS [43, 6], is typically set to 3, meaning the model can output up to three individual separated streams. This configuration constrains the length of the input signal (e.g., input segments are limited to 2.4 seconds in [43] and 3 seconds in [6]). To handle this limitation, chunk-wise decoding is commonly applied, with overlapping regions between adjacent chunks used to align the separated streams. In processing multi-speaker utterances, the CSS-separated results also require corresponding speaker attribution. One approach is to cluster the separated audios to determine the speaker identity for each segment [44]. The NOTSOFAR-1 baseline [6] utilizes the Whisper “large-v3” [45] model to transcribe the CSS-separated results, providing both the text and corresponding word-level timestamps. These timestamped segments are then assigned to specific speakers based on spectral clustering (SC), namely the “CSS + SC” framework.
@@ III Deep Cascade Fusion of Diarization and Separation Framework
Fig. 2 illustrates the overall DCF-DS framework, which is composed of three main components: sliding window on speaker priors (SWSP), unified diarization and separation (UDS), and speaker index-based stitching (SIS). The first two components are utilized during both training and decoding, while the SIS component is employed exclusively during decoding to generate globally separated streams that are consistent with the speaker identities in the global context. Similar to the NSD-MS2S system, DCF-DS also requires an initial speaker diarization result, which may include more speakers than the number of separated streams produced by DCF-DS (for example, as shown in the Fig. 2, there are 4 global speakers, but only 3 separated output streams). To handle this, the SWSP process converts the original global speaker distribution into a windowed distribution, ensuring that the number of speakers within each window is generally less than or equal to the number of output nodes. This allows the window-based speaker priors to be fed into the UDS component, which performs window-based diarization and separation. Finally, the SIS component uses global indices to stitch together the window-based separation results, yielding the complete global speech separation outputs, as depicted at the bottom of Fig. 2. The details of each module will be introduced in the following sections.
@@ III-A Sliding Window on Speaker Priors
where indicates whether the -th speaker is active at frame , with a value of 1 representing presence and 0 representing absence. and are global indices. denotes the total number of speakers in , and represents the total frame number of the entire utterance. To manage the large number of speakers and the long duration of , we utilize the sliding window on speaker priors (SWSP) module. This module segments the global speaker priors into smaller, window-based segments using a fixed-length sliding window. The window-based segment is defined as:
where denotes the speaker priors within the -th window, indicates the index of speakers within the window, is the starting frame index of the -th window, and is the fixed window frame length. represents the maximum number of speakers within a window.
During the SWSP process, we compact the speaker index within each window. For example, in the second window at the top of Fig. 2, we convert the global indices into the windowed indices , effectively controlling the number of speakers within the window. We also save this correspondence using the mapping function for the -th window, which will be used in the subsequent decoding process. In rare cases during decoding, if the number of speakers within a window exceeds , we select the speakers with the longest duration in the priors. During training, such windows are discarded and not used for training. Conversely, if is greater than the actual number of speakers in the current window, we pad with a zero vector. Note that we use manually annotated global segmentation during training, while in decoding, we employ an auxiliary speaker diarization system to provide the global segmentation priors as in NSD-MS2S. The SWSP approach ensures that the number of separation output nodes remains not too large, which allows each separation node to be sufficiently activated during training. During decoding, the SWSP approach enables a limited number of separation output nodes to handle an arbitrary number of global speakers while preserving crucial long context information in .
@@ III-B Unified Diarization and Separation
Once the windowed prior is obtained, it can be used to perform joint diarization and separation, as illustrated in Fig. 2. This module contains the core neural network component of the DCF-DS framework, with its training process illustrated in Fig. 3. The training process involves two main workflows: the primary joint training process and an optional training process for the MIMO-SE module. In the first part, we input the speaker embeddings , corresponding to the speakers within the -th window, along with the window-based FBANKs , into the sequential joint model for diarization and separation. First, these features are fed into the NSD-MS2S architecture as shown in Fig. 3. The process can be described as:
@@
where represents the probability of speaker presence for speakers across frames within the -th window. After obtaining , we utilize the soft boundary information it provides to enhance the performance of the speech separation module. Assume that the time-domain signal within the -th window is denoted as , we can obtain the STFT magnitude feature as:
@@
We then perform a dimensional extension (DE) by replicates along an additional axis, which is treated as the frequency dimension to ensure alignment with the STFT magnitude feature. This operation can be expressed as:
@@
The added dimension ensures that matches the shape of the STFT magnitude feature, enabling the combined features to be processed by the separation model:
@@
The concatenated features combine spectral information and temporal boundaries into a unified input for the separation model. This integration effectively supplies the model with speaker identity, order, and temporal boundary priors, mitigating the permutation problem [46]. Additionally, comparing Eqs. (11) and (18), we can see that this unified input changes the nature of the separation task: instead of mapping directly from the mixed speech signals, the model now maps from the time-frame distribution of speakers to their time-frequency distribution. This adjustment simplifies the speech separation task compared to the traditional approach. After the separation model, we introduce an optional MIMO-SE module, which directly utilizes the Conformer-based speech separation outputs in Eq. (18) to generate the cleaner T-F mask for each speaker. The MIMO-SE takes the concatenated features of the separation results and the STFT magnitude features as input:
@@
Then, the MIMO-SE module uses the Conformer architecture to process these concatenated features, producing the enhanced T-F mask for each speaker:
@@
The input and output dimensions in Eq. (20) are consistent with those of the Conformer-based speech separation model in Eq. (18), allowing for the reuse of the same Conformer architecture. By comparing Eqs. (11), (18), and (20), we see that MIMO-SE serves as an extension of the separation module, simplifying the task further: rather than mapping from a T-mask to a T-F mask as in Eq. (18), it now maps from a T-F mask to a cleaner T-F mask. The enhancement module aims to refine the quality of the separated speech by utilizing the initial separation results, ultimately producing more robust and cleaner separated streams.
@@
where and are elements of the matrices and , respectively, representing the predicted present probabilities and ground truth labels for speaker at frame within the -th window. For the T-F masks produced by the separation module, we use mean squared error (MSE) loss:
where represents the ground truth T-F mask for speech separation defined as follows:
where is the magnitude of the clean speech for speaker under the -th window. The overall loss for joint training is a weighted combination of the BCE and MSE losses:
where controls the weighting of the BCE loss, allowing us to balance the contributions of the diarization and separation tasks during training. In the overall training process, we optimize both the diarization and the separation components. This is because training solely with the separation loss would reduce the entire joint model to merely a large speech separation model, thereby negating the purpose of joint training. Introducing the diarization loss ensures that the diarization module continues to provide speaker boundary information. Additionally, it prevents the speech separation module from over-fitting to the probabilities predicted by the diarization module (the predicted results usually contain inevitable errors). This is achieved by always updating the diarization model’s parameters with the correct labels during training, thereby maintaining the robustness of the entire system.

### v2

@@ Abstract
We propose a single-channel Deep Cascade Fusion of Diarization and Separation (DCF-DS) framework for back-end speech recognition, combining neural speaker diarization (NSD) and speech separation (SS). First, we sequentially integrate the NSD and SS modules within a joint training framework, enabling the separation module to leverage speaker time boundaries from the diarization module effectively. Then, to complement DCF-DS training, we introduce a window-level decoding scheme that allows the DCF-DS framework to handle the sparse data convergence instability (SDCI) problem. We also explore using an NSD system trained on real datasets to provide more accurate speaker boundaries during decoding. Additionally, we incorporate an optional multi-input multi-output speech enhancement module (MIMO-SE) within the DCF-DS framework, which offers further performance gains. Finally, we enhance diarization results by re-clustering DCF-DS outputs, improving ASR accuracy. By incorporating the DCF-DS method, we achieved first place in the realistic single-channel track of the CHiME-8 NOTSOFAR-1 challenge. We also perform the evaluation on the open LibriCSS dataset, achieving a new state-of-the-art single-channel speech recognition performance.
@@
Multi-speaker speech recognition aims to identify “who spoke what and when” in recordings with unknown scenarios and speakers [1]. While deep learning has driven significant advancements in speech recognition [2, 3], accurate speech recognition in real-world multi-speaker conversational scenarios remains challenging. The difficulties often arise from high overlap rates, background noises and reverberation, unknown speaker numbers, and natural conversation styles [4, 5, 6]. In this context, speaker diarization and speech separation are commonly employed as pre-processing steps. Specifically, speaker diarization segments the audio based on speaker identity [7, 8], and speech separation [9] isolates each speaker’s speech from the mixed audio signals.
@@
Speaker diarization and speech separation are closely related tasks [4], as they both aim to predict the distribution of speakers: speaker diarization predicts the speaker’s presence across time frames, whereas speech separation predicts their distribution across time-frequency (T-F) bins. Consequently, the temporal distribution provided by diarization can serve as a prior for speech separation, helping the system to better distinguish between speakers [10, 11, 12, 13, 4, 14]. Speech separation, in turn, inherently encompasses diarization—ideal separation results reveal the time regions of speakers [15, 16, 17]. Moreover, speech separation can handle overlapping segments, which can enhance speaker diarization accuracy [18, 6]. Due to the strong correlation between speaker diarization and speech separation, numerous kinds of research [19, 20, 21, 18, 6, 15, 16, 17, 10, 11, 12, 13, 4, 14, 22] have been conducted to explore combining these two tasks to enhance speech quality. Based on how the two tasks are combined, these methods can be broadly categorized into multi-task joint training and sequential cascade approaches.
@@
In multi-task joint training approaches [19, 20, 22, 21], the network typically generates two outputs: speaker activity probabilities and separated speech. This structure allows the two tasks to benefit from each other mutually [20]. Speaker diarization results can also be used to post-process the separation outputs, reducing background noises in the separated results [19, 20, 21, 22]. Among these methods, some studies [19, 22] propose a network that jointly performs target speech extraction and voice activity detection (VAD), yielding improvements in both fully and sparsely overlapped speech. Another approach, end-to-end neural speaker diarization and separation (EEND-SS) [20], integrates speaker counting, diarization, and separation in a multi-tasking framework. However, due to the inclusion of the separation task within the joint training framework, these methods are restricted to training on simulated data. Unsupervised training methods, such as PixIT [21], can help mitigate this issue. However, when the number of separated output nodes is limited, clustering methods are still required to assign separated streams to speakers for long recordings.
@@
In sequential cascade framework [18, 6, 15, 16, 17, 10, 11, 12, 13, 4, 14], there is no definitive answer as to whether speech separation or speaker diarization should be performed first. Some approaches [18, 6, 15, 16, 17] choose to perform speech separation first, namely “separation-then-diarization”. For example, the continuous speaker separation (CSS) pipeline [18, 6] first processes overlapping regions in the original audios through speech separation and then applies the separated results to diarization. Some studies [15, 16, 17] explore speaker diarization by performing VAD on the separated outputs, such as speech separation guided diarization (SSGD) [15] and recurrent selective attention network (RSAN) [16, 17]. Some methods [10, 11, 12, 13, 4, 14] begin with speaker diarization and then use the obtained time boundary information to guide speech separation, namely “diarization-then-separation”. This is also a promising direction, as speaker diarization is generally simpler than speech separation, allowing the former to provide a more accurate prior. A classic method is guided source separation (GSS) [10, 11], which uses time boundaries to initialize time-varying mixture weights based on traditional spatial mixture models [23, 24]. However, GSS relies on multi-channel spatial information. The speaker activity driven speech extraction neural network (ADEnet) [25] explores how to use speaker activity boundaries as auxiliary cues for speech extraction. The CSS-AD [26] pipeline performs speech recognition on CSS outputs, followed by diarization on the ASR output, achieving promising results on the LibriCSS dataset [18]. Certain approaches [12, 13] enhance separation performance in real two-speaker scenarios by using speaker diarization to generate adaption training data. Recently, the target-speaker separation (TS-SEP) [4] and speaker separation via neural diarization (SSND) [14] have achieved strong performance in the fields of diarization and separation. TS-SEP method combines speaker diarization and speech separation by extending the output of target-speaker voice activity detection (TS-VAD) [27] from the time domain to the time-frequency domain. SSND uses estimated speaker boundaries to assign speakers to the outputs from a multi-speaker separation model. These two methods demonstrate the great potential of the sequential cascade framework.
@@
From various studies mentioned above, it is apparent that both joint training and sequential cascade approaches have advantages and limitations. Joint training allows simultaneous diarization and separation within a compact and unified framework. However, typical multi-task learning frameworks usually do not fully leverage diarization results to optimize separation results. Additionally, joint training frameworks are generally restricted to simulated data. Conversely, the sequential cascade approach—particularly “diarization-then-separation”—allows for more effective integration of two tasks, where separation can effectively leverage diarization results. However, this approach tends to be with a more cumbersome process and independently optimized components. Therefore, an effective approach is to combine “joint training” and “sequential cascade“, thereby taking advantage of both methods—what we refer to as “deep cascade fusion”. However, due to the differences in diarization and separation tasks, directly integrating them introduces a problem: diarization models can have many output nodes (e.g., eight output nodes in [28]). Directly feeding these outputs into the separation model and aligning the number of separation outputs with those of the diarization module may lead to non-convergence [4]. This issue arises because the separation model usually allocates an output node for each speaker. However, when training on typical conversational data, each node has limited target signals, making it difficult for the neural network to learn speaker-discriminative features. For example, in a natural eight-speaker meeting conversation, there are many regions where only one or two people are speaking, a separation model with eight outputs will have only one or two nodes with effective training targets in those regions. This significantly increases the training difficulty, which we term the sparse data convergence instability (SDCI) problem. This issue was also highlighted when the TS-SEP method was proposed [4], where a two-stage training approach was proposed as the solution. However, two-stage training makes the TS-SEP method unable to fully leverage the boundary information from the diarization module for separation. The SSND addresses this issue by constructing dual embedding sequences, which leads to a complex framework. This approach requires a pre-trained diarization model to extract speaker embeddings, followed by constructing an embedding sequence based on boundary information during both the training and testing phases. Additionally, a scenario-aware differentiated loss [22, 29] has been proposed to handle the sparsely overlapped speech. However, this approach requires defining multiple different loss functions and prior knowledge of each speaker’s VAD information.
@@
To address the above issues, we propose a single-channel deep cascade fusion of diarization and separation (DCF-DS) framework for back-end multi-speaker speech recognition. The main contributions of our study are as follows:
@@
In DCF-DS, a joint training framework sequentially combining neural speaker diarization and separation is designed. Compared to TS-SEP, DCF-DS effectively leverages temporal boundaries from the diarization, reducing the difficulty of separation. Compared to SSND, DCF-DS avoids complex pre-processes and eliminates dependence on spatial information. To complement DCF-DS training, we design a window-level decoding scheme to handle the SDCI problem. This decoding scheme allows the separation module to produce fewer outputs than the global number of speakers in an utterance.
@@ II-A Neural Speaker Diarization
where represents the input acoustic features. In NSD-MS2S, we use log-Mel filter-bank features (FBANKs). denotes the -dimensional feature vector at frame , where represents the total frame number. represents the speaker embedding of the -th speaker, and denotes the number of speakers. is the speaker mask matrix, which has elements that indicate whether speaker is present at frame , taking a value of 1 if present and 0 if not. The speaker mask matrix is obtained through manual annotations during training and an initialized diarization system (e.g., spectral clustering) during testing. The NSD-MS2S transforms the inputs into deep features through the following process:
where denotes the convolutional layers, is the number of output channels. is the feature dimension after downsampling. represents the memory-aware multi-speaker embedding (MA-MSE) module proposed in [42], which can provide dynamically refined speaker embeddings using the speaker embedding basis matrix . Next, NSD-MS2S passes these features through an encoder-decoder architecture to generate the final output:
@@ II-B Continuous Speech Separation
where denotes the T-F mask for speaker . denotes the number of outputs. In traditional CSS [43, 6], is typically set to 3, meaning the model can output up to three separated streams. This configuration constrains the length of the input signals (e.g., input segments are limited to 2.4 seconds in [43] and 3 seconds in [6]). To handle this limitation, chunk-wise decoding is commonly applied, with overlapping regions between adjacent chunks used to align the separated streams. In processing multi-speaker utterances, the CSS-separated results also require corresponding speaker attribution. One approach is to cluster the separated audios to determine the speaker identity for each segment [44]. The NOTSOFAR-1 baseline [6] utilizes the Whisper “large-v3” [45] model to transcribe the CSS-separated results, providing both the text and word-level timestamps. These timestamped segments are then assigned to specific speakers based on spectral clustering (SC), namely the “CSS + SC” framework.
@@ III Deep Cascade Fusion of Diarization and Separation Framework
Fig. 2 illustrates the overall DCF-DS framework, which is composed of three main components: sliding window on speaker priors (SWSP), unified diarization and separation (UDS), and speaker index-based stitching (SIS). The first two components are utilized during both training and decoding, while the SIS component is employed exclusively during decoding to generate globally separated streams that are consistent with the speaker identities in the global context. Similar to the NSD-MS2S system, DCF-DS also requires an initial speaker diarization result, which may include more speakers than the number of separated outputs in the DCF-DS (for example, as shown in the Fig. 2, there are 4 global speakers, but only 3 separated output streams). To handle this, the SWSP process converts the original global speaker distribution into a windowed distribution, ensuring that the number of speakers within each window is generally less than or equal to the number of output nodes. This allows the window-based speaker priors to be fed into the UDS component, which performs window-based diarization and separation. Finally, the SIS component uses global indices to stitch together the window-based separation results, yielding the complete global speech separation outputs, as depicted at the bottom of Fig. 2. The details of each module will be described in the following sections.
@@ III-A Sliding Window on Speaker Priors
where indicates whether the -th speaker is active at frame , with a value of 1 representing presence and 0 representing absence. and are global indices. denotes the total number of speakers in . represents the total frame number of the entire utterance. To manage the large number of speakers and the long duration of , we utilize the sliding window on speaker priors (SWSP) module. This module segments the global speaker priors into smaller, window-based segments using a fixed-length sliding window. The window-based prior is defined as:
where denotes the speaker priors within the -th window. indicates the speaker index within the window. is the starting frame index of the -th window. is the fixed window frame length. represents the maximum speaker number within a window.
During the SWSP process, we compact the speaker index within each window. For example, in the second window at the top of Fig. 2, we convert the global indices into the windowed indices , effectively controlling the number of speakers within the window. We also save this correspondence using the mapping function for the -th window, which will be used in the subsequent decoding process. In rare cases during decoding, if the number of speakers within a window exceeds , we select the speakers with the longest duration in the priors. During training, such windows are discarded and not used for training. Conversely, if is greater than the number of speakers in the current window, we apply the zero-padding. Note that we use manually annotated global diarization priors during training, while in decoding, we employ an auxiliary speaker diarization system to provide the global segmentation priors as in NSD-MS2S. The SWSP approach ensures that the number of separation output nodes remains not too large, which allows each separation node to be sufficiently activated during training. During decoding, the SWSP approach enables a limited number of separation output nodes to handle an arbitrary number of global speakers while preserving crucial long context information in .
@@ III-B Unified Diarization and Separation
Once the windowed prior is obtained, it can be used to perform joint diarization and separation, as illustrated in Fig. 2. This module contains the core neural network component of the DCF-DS framework, with its training process illustrated in Fig. 3. The training process involves two main workflows: the primary joint training process and an optional MIMO-SE training process. In the first part, we input the speaker embeddings (corresponding to the speakers within the -th window) and the window-based FBANKs into the sequential joint model. First, these features are fed into the NSD-MS2S architecture as shown in Fig. 3. The process can be described as:
@@
where represents the probability of speaker presence for speakers across frames within the -th window. After obtaining , we utilize the time boundary information it provides to enhance the performance of the speech separation module. Assume that the time-domain signal within the -th window is denoted as , we can obtain the STFT magnitude feature as:
@@
We then perform a dimensional extension (DE) by replicating along an additional axis, which is treated as the frequency dimension. This operation can be expressed as:
@@
The added dimension ensures that each speaker’s time mask in matches the shape of the STFT magnitude feature as shown in Fig. 2, enabling the combined features to be processed by the separation model:
@@
The concatenated features combine spectral information and temporal boundary information into a unified input. This integration effectively supplies the model with speaker identity, order, and temporal boundary priors, mitigating the permutation problem [46]. Additionally, comparing Eqs. (11) and (18), we can see that this unified input changes the nature of the separation task: instead of mapping directly from the mixed speech signals, the model now maps from the time-frame distribution of speakers to their time-frequency distribution. This adjustment simplifies the speech separation task compared to the traditional approach. After the separation model, we introduce an optional MIMO-SE module, which directly utilizes the Conformer-based speech separation outputs in Eq. (18) to generate the cleaner T-F mask for each speaker. The MIMO-SE takes the combination of the separation results and the STFT magnitude features as input:
@@
where . The MIMO-SE module uses the Conformer architecture to process these concatenated features, producing the enhanced T-F mask for each speaker:
@@
where . The input and output dimensions in Eq. (20) are consistent with those of the Conformer-based speech separation model in Eq. (18), allowing for the reuse of the same Conformer architecture. By comparing Eqs. (11), (18), and (20), we can see that MIMO-SE serves as an extension of the separation module, simplifying the task further: rather than mapping from a T-mask to a T-F mask as in Eq. (18), it now maps from a T-F mask to a cleaner T-F mask. The enhancement module aims to refine the quality of the separated speech by utilizing the initial separation results, ultimately producing more robust and cleaner separated streams.
@@
where and are elements of the matrices and , respectively, representing the predicted present probabilities and ground truth labels for speaker at frame within the -th window. For the T-F masks estimated by the separation module, we use mean absolute error (MAE) loss:
where represents the ground truth T-F mask for speech separation defined as:
where is the magnitude of the clean speech for speaker under the -th window. The overall loss for joint training is a weighted combination of the two losses:
where controls the weight of the BCE loss, allowing us to balance the contributions of the diarization and separation tasks during training. In the overall training process, we optimize both the diarization and the separation components. This is because training solely with the separation loss would reduce the entire joint model to merely a large speech separation model, which defeats the purpose of joint training. Introducing the diarization loss ensures that the diarization module continues to provide time boundary information. Additionally, the predicted probabilities of the diarization module may contain unavoidable errors. The joint training process prevents the speech separation module from over-fitting to these predictions, which constantly updates the diarization model’s parameters using the correct labels and maintains the robustness of the entire system.